# ACEGEN: Reinforcement learning of generative chemical agents for drug discovery

### v1

@@ Abstract
In recent years, reinforcement learning (RL) has emerged as a valuable tool in drug design, offering the potential to propose and optimize molecules with desired properties. However, striking a balance between capability, flexibility, and reliability remains challenging due to the complexity of advanced RL algorithms and the significant reliance on specialized code. In this work, we introduce ACEGEN, a comprehensive and streamlined toolkit tailored for generative drug design, built using TorchRL, a modern decision-making library that offers efficient and thoroughly tested reusable components. ACEGEN provides a robust, flexible, and efficient platform for molecular design. We validate its effectiveness by benchmarking it across various algorithms and conducting multiple drug discovery case studies. ACEGEN is accessible at https://github.com/acellera/acegen-open.
@@ 1 Introduction
Drug design is a complex process that involves the identification of biomolecules with multiple optimal properties, such as potency, selectivity, bioavailability, and toxicity. In recent years, a diversity of generative model solutions have been proposed as a promising approach to partially automate the process of proposing new molecules in the design-make-test-analysis cycle (Elton et al., 2019; Wang et al., 2022). These models typically employ machine learning algorithms to generate molecular candidates, but it remains challenging to efficiently search the vast chemical space to identify molecules with optimal properties (Brown et al., 2019), which is so large that it is not practically enumerable in a naive manner.
@@
Reinforcement learning (RL) has emerged as a possible solution to explore this chemical space (Olivecrona et al., 2017; Popova et al., 2018) with an increasing focus on efficiency (Thomas et al., 2022; Guo et al., 2022). RL (Sutton & Barto, 2018) consists of a family of machine learning algorithms that use feedback as a learning signal. Through a trial and error process, RL algorithms can adapt their search strategy to the specific characteristics of compounds associated with desirable properties, and guide the generative models to propose novel molecules with similar characteristics. These methods use a customized reward function that assigns values to molecules, tailored specifically for the application or research question being addressed. To navigate the molecular space and find the most highly rewarded molecules, RL generative agents alternate between two phases: a data collection phase and a learning phase. During the data collection, molecules are generated by the generative model and scored based on the reward function. In the agent learning phase, the collected data is utilized to enhance the agent’s ability to generate interesting and relevant compounds. By leveraging this feedback loop, the generative model can be continually improved to generate compounds with increasingly desirable properties. This is particularly interesting when the available chemical space cannot be enumerated and filtered using the scoring function, either because the scoring functions are very slow or because the search space is very large.
@@
Currently, available implementations for drug discovery with RL rely heavily on custom code (Blaschke et al., 2020; Gao et al., 2022), favoring redundancy, complexity, limited efficiency and flexibility to incorporate a diverse set of solutions. However, the RL community already provides solutions to address these challenges. TorchRL, a comprehensive RL library (Bou et al., 2023), offers well-tested independent RL state-of-the-art components. In this work, we adopt TorchRL’s components as building blocks to assemble efficient and reliable drug discovery agents - a practice already successfully applied in diverse domains like drone control (Xu et al., 2024) and combinatorial optimization (Berto et al., 2023). This approach also fosters algorithmic research by enabling the encapsulation of new ideas within new components, which can seamlessly integrate with existing ones. TorchRL operates within the PyTorch ecosystem (Paszke et al., 2019), ensuring not only high-quality standards but also maintenance over time and future development.
@@
To showcase the advantages of ACEGEN, we implement and evaluate well-known language-based algorithms for drug design. Language models can learn complex patterns in text and generate novel sequences, including molecular structures (Segler et al., 2018; Gómez-Bombarelli et al., 2018). Studies have focused on utilizing Simplified Molecular Input Line Entry System (SMILES) (Weininger, 1988), a string notation system used to represent chemical structures in a compact and standardized manner. Moreover, several prior studies have evidenced the benefits of combining generative language models and RL (Olivecrona et al., 2017; Fialková et al., 2021; Guo et al., 2022). Therefore, we re-implement three REINFORCE-based algorithms, REINFORCEWilliams (1992), REINVENT (Olivecrona et al., 2017) and AHC (Thomas et al., 2022). We also adapt general RL algorithms such as A2C (Mnih et al., 2016) and PPO (Schulman et al., 2017) to our problem setting. The REINVENT and the AHC methods use some level of experience replay, therefore we also add experience replay to REINFORCE. For this same reason, we also test PPOD (Libardi et al., 2021), a PPO-based algorithm adapted for experience replay.
@@
To demonstrate the possible use cases and applications of ACEGEN, we have benchmarked these RL algorithms on a sample efficiency benchmark while considering chemistry, conducted an ablation study to better understand the components of a specific RL algorithm, tested a more challenging drug design relevant objective, and lastly presented RL for molecular constrained generative design.
ACEGEN code is accessible open source under the MIT license at https://github.com/acellera/acegen-open to facilitate uptake by the community.
@@ 2.1 Reinforcement learning setting
Reinforcement learning tasks are formalized as Markov Decision Processes (MDPs) (Sutton & Barto, 2018). Formally, a MDP is described by a quintuple, denoted as $\langle S,A,R,P,\rho_{0}\rangle$. Here, $S$ represents the set of all possible states in the problem space, $A$ denotes the set of valid actions available to the agent and $R:S\times A\times S\to\mathbb{R}$ defines the reward function, which assigns numerical values to transitions from one state to another based on the action taken. The function $P:S\times A\to\mathcal{P}(S)$ indicates the transition probability, with $P(s_{t+1}|s,a)$ representing the probability of transitioning to state $s_{t+1}$ from the current state $s$ under action $a$. Lastly, $\rho_{0}$ signifies the initial state distribution.
@@
This approach naturally extends to the problem of sequentially designing molecules. In this context, a parameterized reinforcement learning (RL) policy, denoted as $\pi_{\theta}$, guides molecule design by successively processing the current state representing a partially built molecule and selecting actions $a_{t}$ to make changes to the molecule’s structure at each step until reaching a terminal state. The reward function evaluates the quality of the molecule, assigning a scalar value either at each generative step or at the end of the process when a complete compound is produced. Within this framework, $\pi_{\theta}(a_{t}|s_{t})$ denotes the probability of the policy function parameterized by $\theta$ taking action $a_{t}$ in state $s_{t}$. $\tau$ represents a complete sequence of actions needed to construct a molecule, often referred to as an episode. $P(\tau|\theta)$ indicates the probability of a trajectory $\tau$ given the policy parameters $\theta$. Furthermore, $R(\tau)$ represents the cumulative sum of rewards over the trajectory $\tau$.
@@
The goal of RL algorithms is to optimize the parameters of the policy $\pi_{\theta}$ to maximize $R(\tau)$. Different methods within the family of policy gradient methods (Sutton & Barto, 2018), such as REINFORCE (Williams, 1992), A2C (Advantage Actor-Critic) (Mnih et al., 2016), and PPO (Proximal Policy Optimization) (Schulman et al., 2017), are commonly used for this task. Additionally, various techniques can aid in training. Reward shaping involves modifying the reward function to provide additional feedback to the agent during training, potentially speeding up learning by providing more informative signals. Experience replay involves storing past experiences in a buffer and randomly sampling from it during training, which can improve sample efficiency and stabilize learning. Moreover, incorporating penalty terms, such as a Kullback-Leibler (KL) divergence loss term, can be utilized to encourage the agent to stay close to a reference or prior policy, helping to maintain stability and control the exploration-exploitation trade-off during training. Finally, ranking and selecting only the best K molecules in each given batch can improve efficiency. ACEGEN integrates all the aforementioned algorithms and additional techniques to form a comprehensive suite of state-of-the-art methods for molecular generation.
@@ 2.2 Chemical language generative models
One possible approach to sequentially designing molecules is through string representations. Molecular string representations(Wiswesser, 1985) convert molecular graphs into strings and vice versa, allowing molecules to be encoded using specific syntactical rules. The task of molecular generation can then be formulated as a natural language processing (NLP)(Öztürk et al., 2020) problem. These have also been named Chemical Language Models (CLMs) (Grisoni, 2023). In this context, the resulting action space is discrete and each action is denoted also as a token. These actions are integrated by concatenating them to form a joint string representation, producing the next state. Each episode begins with a single special start token and can last for a varying number of steps. The episode ends when the agent chooses another special token called the stop token.
@@
ACEGEN currently provides an environment for language model experimentation including SMILES (Weininger, 1988), DeepSMILES (O’Boyle & Dalke, 2018), SELFIES (Krenn et al., 2020), AtomInSmiles (Ucak et al., 2023) and SAFE (Noutahi et al., 2024) grammars, complemented by a user-friendly vocabulary class.  Appendix A illustrates how the vocabulary and environment can be easily created and utilized for data generation.
@@
In practice, CLMs are first trained unsupervised on a bulk of unlabeled data, aimed at learning to generate valid molecules, also referred to as pre-training, which is achieved through the application of the teacher enforcing method. (Lamb et al., 2016). The pre-trained CLM is the starting policy that can be further trained to optimize a specific objective with RL, similar to what happens in large language models. Moreover, this prior policy can be used as an anchor point from which the new RL policy should not deviate excessively.
@@
ACEGEN currently provides pre-trained models for several architectures. More specifically, we provide GRU (Cho et al., 2014) policies pre-trained on two datasets: ChEMBL (Gaulton et al., 2012) and ZINC Gómez-Bombarelli et al. (2018). We also provide weights for an LSTM (Hochreiter & Schmidhuber, 1997) policy pre-trained on ChEMBL (Gaulton et al., 2012) and for a GPT2 (Radford et al., 2019) pre-trained on the REAL Database REAL lead-like compounds (Shivanyuk et al., 2007).
@@
While our repository currently provides methods to create LSTM (Hochreiter & Schmidhuber, 1997), GRU (Cho et al., 2014), GPT2 Radford et al. (2019) , Llama2 (Touvron et al., 2023) and Mamba (Gu & Dao, 2023) policies, practitioners can similarly include any other architectures of their choice taking advantage of current and future developments in NLP. Users can pre-train their policies with external code and integrate them without any need to modify the internal scripts of the library. Our repository contains a step-by-step tutorial that explains how to do the integration.
@@
Finally, ACEGEN also provides its own pre-training script to train language models, either from ACEGEN or custom, to generate valid molecules through the teacher enforcing method (Lamb et al., 2016). Our code is engineered to adapt to the available computational resources, whether a single GPU or a distributed setup spanning multiple machines and GPUs. This adaptability allows us to efficiently train on datasets of large size.
@@ 2.3 Scoring and Evaluation of Molecules
Scoring functions must reflect real-world drug design scenarios and be flexible enough to apply to a range of drug design challenges. Often, overly simplistic objective functions are used for optimization, such as penalized logP Jin et al. (2018), or complex solutions tailored to one particular generative model Blaschke et al. (2020).
ACEGEN allows the integration of custom scoring functions by providing the flexibility to define them as Python methods that accept strings and return numerical values. Alternatively, ACEGEN is optionally compatible with the MolScore Thomas et al. (2023) library, which offers a broad range of drug design relevant scoring functions, diversity filters, support for curriculum learning and also contains a benchmarking mode including MolOpt (Gao et al., 2022), GuacaMol (Brown et al., 2019), and others.  Appendix B showcases how both scoring options can be accommodated within our codebase. Additionally, the library offers a detailed tutorial, guiding users through the process of implementing and incorporating custom scoring functions seamlessly into the entire workflow to train agents with them.
@@ 2.4 RL Agents Training
ACEGEN provides training for RL agents utilizing the methods: REINFORCE (Williams, 1992), REINVENT Blaschke et al. (2020), AHC Thomas et al. (2022), A2C Mnih et al. (2016) and PPO Schulman et al. (2017) and an adaptation of the PPOD Libardi et al. (2021) algorithm. We deviate from the exact implementation of PPOD by omitting the use of an initial expert demonstration, refraining from replaying episodes with high-value predictions (only episodes with high rewards are replayed) and employing a fixed amount of replay data per batch. All methods are fully configurable to allow the use of custom scoring functions and models beyond those already provided by the library.
REIVENT and AHC can be considered extensions of the REINFORCE algorithm. More specifically, REINVENT incorporates reward shaping, experience replay, and a penalty for high-likelihood sequences into the basic REINFORCE framework. Additionally, AHC ranks and selects the top K molecules in each batch of data. We find the reward shaping not to be compatible with advantage-based methods and the high-likelihood penalty term detrimental, therefore for A2C, PPO and PPOD we introduce instead a term to the loss function based on the Kullback-Leibler divergence (KL) between the actor policy and a prior policy. This term serves to penalize the policy for deviating too much from a prior, similar in concept to reward shaping in REINVENT and AHC. The incorporation of KL constraints is a common practice in research papers aligning language models with custom reward functions, as seen in examples utilizing human feedback Menick et al. (2022); Ouyang et al. (2022); Bai et al. (2022).
@@ 2.5 De-novo, decorative and fragment-linking generation
In drug discovery pipelines, generating molecules from scratch may not always suffice. To meet diverse requirements, ACEGEN offers multiple sampling modes. PromptSMILES (Thomas et al., 2024) is a simple method enabling constrained molecule generation using models pre-trained solely with teacher enforcement on full SMILES, making it possible to generate molecules while adhering to specific chemical sub-structures. Specifically, in addition to de-novo generation, ACEGEN scripts allow for easy configuration of scaffold decoration and fragment-linking molecule generation modes. Constrained sampling modifies only the data collection behavior, operating independently of all other agent components, in line with the TorchRL philosophy. Tutorials on performing constrained generation are provided within the repository.

### v2

@@ Abstract
In recent years, reinforcement learning (RL) has emerged as a valuable tool in drug design, offering the potential to propose and optimize molecules with desired properties. However, striking a balance between capabilities, flexibility, reliability, and efficiency remains challenging due to the complexity of advanced RL algorithms and the significant reliance on specialized code. In this work, we introduce ACEGEN, a comprehensive and streamlined toolkit tailored for generative drug design, built using TorchRL, a modern RL library that offers thoroughly tested reusable components. We validate ACEGEN by benchmarking against other published generative modeling algorithms and show comparable or improved performance. We also show examples of ACEGEN applied in multiple drug discovery case studies. ACEGEN is accessible at https://github.com/acellera/acegen-open and available for use under the MIT license.
[…]
@@ 1 Introduction
Drug design is a complex process that involves the identification of biomolecules that have an optimal balance of multiple properties, such as potency, selectivity, bioavailability, and toxicity. In recent years, a diversity of generative modeling solutions have been proposed as a promising approach to partially automate the process of proposing new molecules that simultaneously improve multiple desired properties in the design-make-test-analysis cycle 1, 2. These models typically employ machine learning algorithms to generate molecular candidates, but it remains challenging to efficiently search the vast chemical space to identify molecules with optimal properties 3, which is so large that it is not practically enumerable in a naive manner.
@@
Reinforcement learning (RL) has emerged as a possible solution to explore this chemical space 4, 5 with an increasing focus on efficiency 6, 7. RL 8 consists of a family of machine learning algorithms that use feedback as a learning signal. Through a trial and error process, RL algorithms can adapt their search strategy to the specific characteristics of compounds associated with desirable properties, and guide the generative models to propose novel molecules with similar characteristics. These methods use a customized reward function that assigns values to molecules, tailored specifically for the application or research question being addressed. To navigate the molecular space and find the most highly rewarded molecules, RL generative agents alternate between two phases: a data collection phase and a learning phase. During the data collection, molecules are generated by the generative model and scored based on the reward function. In the agent learning phase, the collected data is utilized to enhance the agent’s ability to generate interesting and relevant compounds. By leveraging this feedback loop, the generative model can be continually improved to generate compounds with increasingly desirable properties. This is particularly interesting when the available chemical space cannot be enumerated and filtered using the scoring function, either because the scoring functions are very slow or because the search space is very large.
@@
Currently, available implementations for drug discovery using reinforcement learning (RL) rely heavily on custom code 9, 10. This approach tends to favor redundancy, complexity, and limited efficiency, making it difficult to incorporate a diverse set of solutions. However, the RL community has already developed solutions to address these challenges. TorchRL, a comprehensive RL library 11, offers well-tested, independent state-of-the-art RL components. In this work, we adopt TorchRL’s components as building blocks to assemble efficient and reliable drug discovery agents—a practice already successfully applied in diverse domains such as drone control 12 and combinatorial optimization 13. This approach also fosters algorithmic research by enabling the encapsulation of new ideas within new components, which can seamlessly integrate with existing ones. TorchRL operates within the PyTorch ecosystem 14, ensuring not only high-quality standards but also maintenance over time and future development.
@@
To showcase the advantages of ACEGEN, we implement and evaluate well-known language-based algorithms for drug design. Language models can learn complex patterns in text and generate novel sequences, including molecular structures 15, 16. Studies have focused on utilizing the Simplified Molecular Input Line Entry System (SMILES) 17, a string notation system used to represent chemical structures in a compact and standardized manner. Moreover, several prior studies have demonstrated the benefits of combining generative language models and RL 4, 18, 7. Therefore, we re-implement three REINFORCE-based algorithms: REINFORCE 19, REINVENT 4, and AHC 6. Additionally, we adapt general RL algorithms such as A2C 20 and PPO 21 to our problem setting. The REINVENT and AHC methods utilize some level of experience replay; therefore, we also incorporate experience replay into REINFORCE. Similarly, we test PPOD 22, a PPO-based algorithm adapted for experience replay.
@@
To demonstrate the possible use cases and applications of ACEGEN, we have benchmarked these RL algorithms on a sample efficiency benchmark while considering chemistry, conducted an ablation study to better understand the components of a specific RL algorithm, tested a more challenging drug design relevant objective, and lastly presented RL for molecular constrained generative design. ACEGEN code is accessible open source under the MIT license at https://github.com/acellera/acegen-open to facilitate uptake by the community.
@@ 2.1 Reinforcement learning setting
Reinforcement learning tasks are formalized as Markov Decision Processes (MDPs) 8. Formally, an MDP is described by a quintuple, denoted as $\langle S,A,R,P,\rho_{0}\rangle$. Here, $S$ represents the set of all possible states in the problem space, $A$ denotes the set of valid actions available to the agent, and $R:S\times A\times S\to\mathbb{R}$ defines the reward function, which assigns numerical values to transitions from one state to another based on the action taken. The function $P:S\times A\to\mathcal{P}(S)$ indicates the transition probability, where $P(s_{t+1}|s,a)$ represents the probability of transitioning to state $s_{t+1}$ from the current state $s$ under action $a$. Lastly, $\rho_{0}$ signifies the initial state distribution.
@@
This approach naturally extends to the problem of sequentially designing molecules. In this context, a parameterized reinforcement learning (RL) policy, denoted as $\pi_{\theta}$, guides molecule design by successively processing the current state representing a partially built molecule and selecting actions $a_{t}$ to make changes to the molecule’s structure at each step until reaching a terminal state. The reward function evaluates the quality of the molecule, assigning a scalar value either at each generative step or at the end of the process when a complete compound is produced. Within this framework, $\pi_{\theta}(a_{t}|s_{t})$ denotes the probability of the policy function, parameterized by $\theta$, taking action $a_{t}$ in state $s_{t}$. The sequence of actions needed to construct a molecule, often referred to as an episode, is represented by $\tau$. $P(\tau|\theta)$ indicates the probability of a trajectory $\tau$ given the policy parameters $\theta$. Furthermore, $R(\tau)$ represents the cumulative sum of rewards over the trajectory $\tau$.
The goal of RL algorithms is to optimize the parameters of the policy $\pi_{\theta}$ to maximize $R(\tau)$. Different methods within the family of policy gradient methods 8, such as REINFORCE 19, A2C (Advantage Actor-Critic) 20, and PPO (Proximal Policy Optimization) 21, are commonly used for this task.
@@
Additionally, various techniques can aid in training. Reward shaping involves modifying the reward function to provide additional feedback to the agent during training, potentially speeding up learning by providing more informative signals. Experience replay involves storing past experiences in a buffer and randomly sampling from it during training, which can improve sample efficiency and stabilize learning. Moreover, incorporating penalty terms, such as a Kullback-Leibler (KL) divergence loss term, can be utilized to encourage the agent to stay close to a reference or prior policy, helping to maintain stability and control the exploration-exploitation trade-off during training. Finally, ranking and selecting only the best K molecules in each given batch can improve efficiency. ACEGEN integrates all the aforementioned algorithms and additional techniques to form a comprehensive suite of state-of-the-art methods for molecular generation.
@@ 2.2 Chemical language generative models
One possible approach to sequentially designing molecules is through string representations. Molecular string representations 23 convert molecular graphs into strings and vice versa, allowing molecules to be encoded using specific syntactical rules. This approach formulates the task of molecular generation as a natural language processing (NLP) 24 problem, and the policy models used to sequentially generate molecules represented as strings are called Chemical Language Models (CLMs) 25. In this context, the resulting action space is discrete, with each action denoted as a token. These actions are integrated by concatenating them to form a joint string representation, producing the next state. Each episode begins with a single special start token and can last for a varying number of steps. The episode ends when the agent chooses another special token called the stop token.
@@
ACEGEN currently provides an environment for language model experimentation including SMILES 17, DeepSMILES 26, SELFIES 27, AtomInSmiles 28 and SAFE 29 grammars, complemented by a user-friendly vocabulary class.  Appendix A A in the appendix illustrates how the vocabulary and environment can be easily created and utilized for data generation in ACEGEN.
@@
In practice, CLMs are first trained unsupervised on a bulk of unlabeled data, aimed at learning to generate valid molecules, also referred to as pre-training, which is achieved through the application of the teacher enforcing method. 30. The pre-trained CLM is the starting policy that can be further trained to optimize a specific objective with RL, similar to what happens in large language models. Moreover, this prior policy can be used as an anchor point from which the new RL policy should not deviate excessively.
@@
ACEGEN currently provides pre-trained models for several architectures. More specifically, we provide GRU 31 policies pre-trained on two datasets: ChEMBL 32 and ZINC 16. We also provide weights for an LSTM 33 policy pre-trained on ChEMBL 32 and for a GPT2 34 pre-trained on the REAL Database REAL lead-like compounds 35.
@@
While our repository currently provides methods to create LSTM 33, GRU 31, GPT2 34 , Llama2 36 and Mamba 37 policies, practitioners can similarly include any other architectures of their choice taking advantage of current and future developments in NLP. Users can pre-train their policies with external code and integrate them without any need to modify the internal scripts of the library. Our repository contains a step-by-step tutorial that explains how to do the integration.
@@
Finally, ACEGEN also provides its own pre-training script to train language models, either from ACEGEN or custom, to generate valid molecules through the teacher enforcing method 30. Our code is engineered to adapt to the available computational resources, whether a single GPU or a distributed setup spanning multiple machines and GPUs. This adaptability allows us to efficiently train on datasets of large size.
@@ 2.3 Scoring and Evaluation of Molecules
Scoring functions must reflect real-world drug design scenarios and be flexible enough to apply to a range of drug design challenges. Often, overly simplistic objective functions are used for optimization, such as penalized logP 38, or complex solutions tailored to one particular generative model 9. ACEGEN allows the integration of custom scoring functions by providing the flexibility to define them as Python methods that accept strings and return numerical values. Alternatively, ACEGEN is optionally compatible with the MolScore 39 library, which offers a broad range of drug design relevant scoring functions, diversity filters, support for curriculum learning and also contains a benchmarking mode including MolOpt 10, GuacaMol 3, and others.  Appendix B B in the appendix showcases how both scoring options can be accommodated within our codebase. Additionally, the library offers a detailed tutorial, guiding users through the process of implementing and incorporating custom scoring functions seamlessly into the entire workflow to train agents with them.
@@ 2.4 RL Agents Training
ACEGEN provides training for RL agents utilizing the following methods: REINFORCE 19, REINVENT 9, AHC 6, A2C 20, and PPO 21, as well as an adapted version of the PPOD algorithm 22. We deviate from the exact implementation of PPOD by omitting the use of an initial expert demonstration, replaying only episodes with high rewards (instead of episodes with high-value predictions), and employing a fixed amount of replay data per batch. All methods are fully configurable, allowing the use of custom scoring functions and models beyond those already provided by the library.
REINVENT and AHC can be considered extensions of the REINFORCE algorithm. Specifically, REINVENT incorporates reward shaping, experience replay, and a penalty for high-likelihood sequences into the basic REINFORCE framework. Additionally, AHC ranks and selects the top K molecules in each batch of data. We find that using the same reward shaping is not compatible with advantage-based methods, and the high-likelihood penalty term is detrimental. Therefore, for A2C, PPO, and PPOD, we introduce a term to the loss function based on the Kullback-Leibler divergence (KL) between the actor policy and a prior policy. This term serves to penalize the policy for deviating too much from a prior, similar in concept to reward shaping in REINVENT and AHC. Incorporating KL constraints is a common practice in research papers aligning language models with custom reward functions, as seen in examples utilizing human feedback 40, 41, 42.
@@ 2.5 De-novo, decorative and fragment-linking generation
In drug discovery pipelines, generating molecules from scratch may not always suffice. To meet diverse requirements, ACEGEN offers multiple sampling modes. PromptSMILES 43 is a simple method enabling constrained molecule generation using models pre-trained solely with teacher enforcement on full SMILES, making it possible to generate molecules while adhering to specific chemical sub-structures. Specifically, in addition to de-novo generation, ACEGEN scripts allow for easy configuration of scaffold decoration and fragment-linking molecule generation modes. Constrained sampling modifies only the data collection behavior, operating independently of all other agent components, in line with the TorchRL philosophy. Tutorials on performing constrained generation are provided within the repository.